OpenStudy (anonymous):

find the length of the curve using integration: 3y=2(x^2+1)^3/2 from x=0 to x=3 ans. 21. I need solution.. please help. thanks:))

OpenStudy (callisto):

\[2/3 \int\limits_{0}^{3} (x ^{2}+1)^{3/2}dx\] \[=2/3 * 2/5 \left[ (1/2x)(x ^{2}+1)^{5/2} \right]_{0}^{3}\] \[=4/15\left[ (1/6)(3 ^{2}+1)^{5/2} \right] \] \[=2(10)^{5/2}/45 \] not sure for the answer sorry

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

l=int(from 0 to 3 ) {sqrt(1+4x^2(x^2+1))} dx solving this u would get the answer as 21.